# Libra: Leveraging Temporal Images for Biomedical Radiology Analysis

### v1

@@ C.4 Impact of Radiology-Specific Pre-trained Models
Aligning radiology images with textual information is a key challenge in RRG tasks. To demonstrate the benefits of using radiology-specific pre-trained models for more accurate feature representation and improved MLLM performance, we initialised a Libra model with RadDINO, the TAC, and Meditron-7b, conducting the first stage of training, denoted as Libra-1 (This is consistent with the baseline setup of the second ablation study in Section 3). Then we replaced the image encoder and LLM with their general-domain counterparts, DINOv2 and Vicuna-7B-v1.5, respectively. Finally, we replaced both components, which is also referred to as Libra-b.
@@
C.5 Incremental Study
In this section, we conducted an incremental study of our model, mirroring the logical framework we used to construct Libra. We started with an architecture similar to LLaVA, comprising a pre-trained CLIP image encoder, a randomly initialised four-layer MLP as the adapter, and Vicuna-7B-v1.5 as the LLM. We train the adapter on the Findings section generation task to establish an initial baseline.
Gradual enhancements are then introduced, as shown in the Table 12. First, the image encoder is replaced with DINOv2. Building upon this, we integrate the LFE, the prefix module of the TAC connector. Subsequently, we add the TFM, the suffix module, thus completing the full TAC connector. Following this, the image encoder and LLM are progressively replaced with biomedical domain models, RAD-DINO and Meditron, respectively. The dataset used for pretraining in the first stage is then extended. Finally, the model is fine-tuned for downstream tasks to obtain Libra.
@@ C.5 Incremental Study
We observed that with each step of improvement, the overall performance consistently exceeded the previous state, demonstrating the essential role of each component in our optimisation scheme. Notably, in the alignment stage, the model showed the greatest overall improvement after the addition of the TFM, highlighting its capability to capture temporal information, which significantly enhanced performance on the RRA task.
@@
However, after the data expansion optimisation in the first stage, the scores on lexical metrics improved, while clinical metrics exhibited a slight decline. This can be attributed to the introduction of the VQA task, which led the model to perceive more fine-grained grounded information. This is distinct from the report generation task, which focuses on describing the holistic level. Furthermore, this shift also affects the F1 score, as temporal entities are often associated with the presence of specific symptoms. Ultimately, these dropped scores are improved again through the second stage of fine-tuning.
@@
Evaluation of Libra’s Temporal Awareness. Another approach to investigating the model’s ability to capture temporal information is to evaluate it separately within the test split based on the presence or absence of true prior images, as shown in Table 13.
@@
With the addition of the TFM, the model exhibited temporal awareness. It is worth noting that, for the first time, the F1 score of samples with prior images surpassed those without, and this trend persisted through subsequent optimisations. This indicates that the structural enhancements have resulted in a sustained improvement in the model’s temporal perception capabilities. An effective example is in Section 5.
Appendix D Appendix D
D.1 Discussion on Performance with Radiology Foundation Models
| Metric | MedVersa | MAIRA-2 | M4CXR | Libra (%) |
| Lexical: | ||||
| ROUGE-L | – | 38.4 | 28.5 | 36.7 (-4.4%) |
| BLEU-1 | – | 46.5 | 33.9 | 51.3 (10.3%) |
| BLEU-4 | 17.8 | 23.4 | 10.3 | 24.5 (4.7%) |
| METEOR | – | 42.0 | – | 48.9 (16.4%) |
| BERTScore | 49.7 | – | – | 62.5 (25.8%) |
| Clinical: | ||||
| RadGraph-F1 | 28.0 | 34.6 | 21.8 | 32.9 (-4.9%) |
| RG | – | 39.7 | – | 37.6 (-5.3%) |
| RadCliQ() | 2.7 | 2.6 | – | 2.7 (-3.8%) |
| CheXbert vector | 46.4 | 50.6 | – | 46.9 (-7.3%) |
| CheXpert-F1: | ||||
| Micro-F1-14 | – | 58.5 | 60.6 | 55.9 (-7.8%) |
| Macro-F1-14 | – | 42.7 | 40.0 | 40.4 (-5.4%) |
| Micro-F1-5 | – | 58.9 | 61.8 | 60.1 (-2.8%) |
| Macro-F1-5 | – | 51.5 | 49.5 | 53.8 (4.5%) |
As shown in Table 14, these models belong to the category of radiology foundation models. MedVersa (Zhou et al., 2024) and M4CXR (Park et al., 2024) support multiple tasks, including medical report generation, visual grounding, and visual question answering. In contrast, MAIRA-2 (Bannur et al., 2024) specialises in grounded radiology report generation, a novel task requiring image-level localisation for each identifiable finding or symptom. It is worth noting that the test sets differ slightly. All of these models utilise additional radiology information, such as lateral views, prior study reports of CXRs, or both.
D.2 Discussion on Performance with non-LLM-based Models
| Model | Lexical Metrics | Clinical Metrics | |||||||
| B-1 | B-2 | B-3 | B-4 | MTR | R-L | P | R | F1 | |
| R2Gen | 35.3 | 21.8 | 14.5 | 10.3 | 14.2 | 27.0 | 33.3 | 27.3 | 27.6 |
| R2GenCMN | 35.3 | 21.8 | 14.8 | 10.6 | 14.2 | 27.8 | 34.4 | 27.5 | 27.8 |
| M2TR | 37.8 | 23.2 | 15.4 | 10.7 | 14.5 | 27.2 | 24.0 | 42.8 | 30.8 |
| KnowMAT | 36.3 | 22.8 | 15.6 | 11.5 | – | 28.4 | 45.8 | 34.8 | 37.1 |
| CMM-RL | 38.1 | 23.2 | 15.5 | 10.9 | 15.1 | 28.7 | 34.2 | 29.4 | 29.2 |
| CMCA | 36.0 | 22.7 | 15.6 | 11.7 | 14.8 | 28.7 | 44.4 | 29.7 | 35.6 |
| KiUT | 39.3 | 24.3 | 15.9 | 11.3 | 16.0 | 28.5 | 37.1 | 31.8 | 32.1 |
| DCL | – | – | – | 10.9 | 15.0 | 28.4 | 47.1 | 35.2 | 37.3 |
| METrans | 25.0 | 16.9 | 12.4 | 15.2 | – | 29.1 | 36.4 | 30.9 | 31.1 |
| ORGAN | 38.6 | 25.6 | 17.2 | 12.3 | 16.2 | 29.3 | 41.6 | 41.8 | 38.5 |
| COMG | 36.3 | 23.5 | 16.7 | 12.4 | 12.8 | 29.0 | – | – | – |
| MedM2G | 41.2 | 26.9 | 17.9 | 14.2 | – | 30.9 | – | – | – |
| BioViL-T | – | – | – | 9.2 | – | 29.6 | – | – | 17.5 |
| TiBiX | 32.4 | 23.4 | 18.5 | 15.7 | 16.2 | 33.1 | 30.0 | 22.4 | 25.0 |
| RECAP | 42.9 | 26.7 | 17.7 | 12.5 | 16.8 | 28.8 | 38.9 | 44.3 | 39.3 |
| Libra | 51.3 | 38.0 | 30.0 | 24.5 | 48.9 | 36.7 | 59.7 | 52.5 | 55.9 |
To facilitate comparison with non-LLM-based models, we selected evaluation metrics commonly used in these studies. These include BLEU- (Papineni et al., 2002), METEOR (MTR) (Banerjee & Lavie, 2005), and ROUGE-L (R-L) (Lin, 2004). For clinical metrics, we adopted the CheXbert (Irvin et al., 2019a), reporting Precision (P), Recall (R), and F1.
Baseline For performance evaluation, we compare our model with the following baselines: R2Gen (Chen et al., 2021), R2GenCMN (Chen et al., 2021), M2TR (Nooralahzadeh et al., 2021), KnowMAT (Yang et al., 2022), CMM-RL (Qin & Song, 2022), CMCA (Song et al., 2022), KiUT (Huang et al., 2023), DCL (Li et al., 2023b), METrans (Wang et al., 2023b), ORGAN (Hou et al., 2023b), COMG (Gu et al., 2023), MedM2G (Zhan et al., 2024), BioViL-T (Bannur et al., 2023), TiBiX (Sanjeev et al., 2024), RECAP (Hou et al., 2023a).
As the results demonstrate, our model, like other LLM-based models, generally outperforms non-LLM-based models. This is attributed to advancements in LLMs and visual instruction tuning (Liu et al., 2023), enabling multimodal large language models (MLLMs) to excel in the RRG task.
Appendix E Appendix E
E.1 Temporal Alignment Heatmap Analysis and Feature Representation
Current image
RAD-DINO layer 1
RAD-DINO layer 12
LFE
TAC
@@ E.1 Temporal Alignment Heatmap Analysis and Feature Representation
The heatmap in Figure 4 corresponds to the example in Figure 2, where no true prior image was used as a reference. It highlights the clear differences in feature representations across layers of the RAD-DINO image encoder. The shallow layers primarily capture the overall lung structure, while the deeper layers focus on specific disease locations.
@@
After being processed by the Layerwise Feature Extractor (LFE), the image feature representations show that larger symptom regions are assigned higher weights. This indicates that the image representation achieves a more fine-grained level of detail. Following the Temporal Alignment Connector (TAC), Libra integrates the weighted dummy prior image, resulting in a more uniform feature distribution. This reflects temporal information, indicating no significant changes compared to the prior study. This process helps the model provide a smoother image feature representation for subsequent processing by the text decoder (i.e., the LLM).
Current image
Current LFE
TAC-1
Prior LFE
Prior image
Swap order
Current image (Prior)
Current LFE (Prior)
TAC-2
Prior LFE (Current)
Prior image (Current)
@@
The heatmap in Figure 5 corresponds to the example in Figure 3, where a true prior image is provided. As observed, after processing through the LFE, the model captures higher-granularity feature representations in the symptom areas. After being processed by the TAC, the intermediate representation not only integrates these fine-grained features but also incorporates the difference information between the two images, reflecting the temporal information provided by the reference image, as shown in TAC-1 (Figure 5).
@@
When the image order is swapped, treating the true prior image as the current image, the intermediate output from the LFE remains unchanged. Nevertheless, a comparison of the TAC-2 and TAC-1 outputs reveals noticeable differences in the lung feature representations. This demonstrates that the model’s temporal perception is directional and that the TAC module successfully integrates temporal information from different time points. At the same time, it confirms that the LFE focuses solely on extracting image feature representations without encoding temporal aspects.
This result is expected because the residual connections in the TAC ensure that the current image remains the main modality, while the prior image serves as the auxiliary modality. When the order is swapped, the prior image becomes the main modality for subsequent processing. Thus, swapping the image order also changes the content of the generated report, as mentioned in Section 5. The primary difference lies in the reversal of the temporal state of symptoms, such as changing from “improving” to “worsening”.

### v2

@@ H.5 Impact of Radiology-Specific Pre-trained Models on Libra
Aligning radiology images with textual information is a key challenge in RRG tasks. To demonstrate the benefits of using radiology-specific pre-trained models for more accurate feature representation and improved MLLM performance, we initialised a Libra model with RadDINO, the TAC, and Meditron-7b, conducting the first stage of training, denoted as Libra-1 (This is consistent with the baseline setup of the previous ablation study in Sec. 4). Then we replaced the image encoder and LLM with their general-domain counterparts, DINOv2 and Vicuna-7B-v1.5, respectively. Finally, we replaced both components, which is also referred to as Libra-b (in Table 10).
@@
H.6 Incremental Component Analysis
@@
We conducted an incremental study to evaluate the effectiveness of each component in Libra’s architecture. Starting with a baseline model similar to LLaVA—comprising a pre-trained CLIP image encoder, a randomly initialised four-layer MLP adapter, and Vicuna-7B-v1.5 as the LLM—we trained the adapter on the Findings section generation task.
Improvements were introduced incrementally, as summarised in Table 14. First, we replaced the image encoder with DINOv2. Next, we incorporated the LFE (prefix module of TAC) and subsequently added the TFM (suffix module), completing the TAC connector. We then replaced the image encoder and LLM with RAD-DINO and Meditron, respectively. The dataset for the first stage was expanded, and final fine-tuning was conducted for downstream tasks to produce Libra.
@@
With each enhancement, the model’s performance improved, demonstrating the critical role of each component. Notably, the addition of the TFM during the alignment stage provided the most significant improvement, showcasing its ability to capture temporal information, which is essential for the RRG task.
@@
However, data expansion in the first stage led to improved lexical scores but a slight decline in clinical metrics, likely due to the VQA task’s focus on fine-grained grounded information rather than holistic report generation, as mentioned in Sec. 4. This shift also affected the F1 score, as temporal entities are often linked to specific symptoms. These declines were subsequently addressed through second-stage fine-tuning, resulting in overall improved performance.
Evaluation of Libra’s Temporal Awareness
@@
Another approach to investigating the model’s ability to capture temporal information is to evaluate it separately within the test split based on the presence or absence of prior images, as shown in Table 15.
@@
With the addition of the TFM, the model exhibited temporal awareness. It is worth noting that, for the first time, the F1 score of samples with prior images surpassed those without, and this trend persisted through subsequent optimisations. This indicates that the structural enhancements have resulted in a sustained improvement in the model’s temporal perception capabilities. An effective example is in Sec. 5.
Appendix I Heatmap Analysis and Temporal Feature Representation
@@
The heatmap in Figure 4 corresponds to the example in (a) of Figure 3, where no prior image was used as a reference. It illustrates the clear differences in feature representations across layers of the RAD-DINO (Pérez-García et al., 2024) image encoder. The shallow layers primarily capture the overall lung structure, while the deeper layers focus on specific disease regions.
@@
After passing through the Layerwise Feature Extractor (LFE), the image feature representations assign higher weights to larger symptom regions, achieving finer granularity. Following the Temporal Alignment Connector (TAC), the model integrates the weighted dummy prior image, producing a uniform feature distribution that reflects temporal information. This indicates no significant changes compared to the prior study and facilitates smoother image feature representations for downstream text generation by the LLM.
@@
The heatmap in Figure 5 corresponds to the example in (b) of Figure 3, where a prior image is provided. After processing through the LFE, the model captures fine-grained feature representations in symptom areas. When processed by the TAC, these features are integrated with the differences between the two images, effectively reflecting temporal information, as demonstrated in TAC-1 (top) of Figure 5.
@@
When the image order is swapped, treating the prior image as the current image, the LFE output remains unchanged. However, comparing TAC-2 (bottom of Figure 5) and TAC-1 outputs reveals significant differences in lung feature representations. This highlights the model’s directional temporal perception and confirms that the TAC module effectively encodes temporal information from different time points, while the LFE focuses solely on image features without temporal encoding.
This behaviour aligns with the design of the TAC, where residual connections prioritise the current image as the main modality and the prior image as the auxiliary. Swapping the image order changes the main modality, altering the temporal state of symptoms in the generated report, such as reversing descriptions from “improving” to “worsening,” as discussed in Sec. 5.
Appendix J Extended Discussion on Limitations
Handling Multiple Prior Scans
Our current model is designed to process a single prior scan alongside the current scan. While this approach aligns with standard clinical workflows, which typically prioritise the most recent prior study for comparisons, it overlooks scenarios where multiple prior scans could offer a richer temporal perspective. For instance, analysing a sequence of images spanning an extended period could provide deeper insights into gradual disease progression. Future efforts should focus on extending our framework to incorporate multiple prior scans efficiently, enabling a more nuanced understanding of temporal patterns in clinical data.
Temporal Information Beyond Image Comparisons
Currently, our model captures temporal information through paired image comparisons and corresponding textual reports. However, clinical assessments often draw upon a broader context, including historical notes, laboratory results, and other longitudinal patient data. Expanding our approach to integrate these diverse temporal data sources could facilitate a more holistic understanding of disease trajectories and patient history, significantly enhancing clinical applicability.
Sparse Temporal Data Challenges
In cases where prior scans are unavailable or minimally informative (e.g., taken within a short interval), our “dummy prior image” provides a workaround. However, the model’s ability to interpret and generate meaningful outputs under these constraints may still be limited. Future research could focus on synthesising or imputing temporal context to enhance performance under these constraints.
Computational Complexity
The use of temporal alignment mechanisms and multi-layer feature integration increases computational demands, posing challenges for deployment in resource-constrained environments. Future optimisation efforts should focus on reducing computational overhead while maintaining performance.
Generalisability Across Modalities and Datasets
Our study is limited to frontal-view chest X-rays and the MIMIC-CXR dataset (Johnson et al., 2019b). The applicability of our approach to other imaging modalities (e.g., CT, MRI) and datasets (e.g., CheXpert (Irvin et al., 2019), PadChest (Bustos et al., 2020)) remains unexplored. Future studies should assess the model’s generalisability to a broader range of datasets and imaging contexts.
Based on the identified limitations, we outline the following directions:
Develop frameworks for integrating multiple prior scans with dynamic temporal reasoning to better capture longitudinal changes.
Expand the model to incorporate multi-modal imaging and textual data for more comprehensive diagnostic insights.
Investigate the integration of diverse temporal data sources, such as electronic health records (EHRs), to enhance clinical applicability.
These advancements aim to address the current limitations while broadening the applicability of temporal-aware multimodal models in radiology and other clinical domains.